# Inflation with vector fields revisited: heavy entropy perturbations and primordial black holes

### v1

@@ Abstract
We revisit inflation coupled with vector fields employing kinetic coupling in the comoving gauge. It is known that there is a cumulative effect $IN^{2}$ on curvature power spectrum. For a large number of e-foldings $N$, this contribution is so significant that it could violate observational constraints when the ratio of kinetic energy between vector fields and inflaton $I$ is not extremely small. In this paper, we explore the regime where $I\gg 1$, a realm that has not been extensively explored due to the limitations of perturbative methods. We found that the entropy perturbation becomes heavy in this regime and the cumulative effect decays away on super-horizon scales. Consequently, the power spectrum retains its scale invariance in the decoupling limit. By straightforwardly integrating out the heavy modes near horizon-crossing, we derive a low-energy effective field theory describing a massless adiabatic perturbation with an imaginary speed of sound $c_{s}^{2}=-1/3$. Namely, the inflation with vector fields presents a potential mechanism for generating primordial black holes.
@@ 1 Introduction
The study of introducing extra fields during inflation has attracted attention to evade the issues of single-field inflation. In particular, the study of multi-scalar fields exhibiting strongly non-geodesic motion has been extensively examined in recent years [8, 9, 10, 11, 12, 13, 14]. In contrast, our focus is on U(1) gauge fields, commonly known as vector fields, as additional fields during inflation. This choice stems not only from the ubiquity of gauge fields as essential components of the standard model (SM) and beyond but also from their rich phenomenological implications in cosmology [15, 16, 17, 18, 19]. The breaking of the conformal coupling between vector fields and gravity is necessary to generate sufficiently strong vector fields [19]. Analogous to the non-geodesic motion of scalar fields, vector fields should exhibit background dynamics and participate alongside scalar fields in driving inflation. The inflationary phase undergoes a transition to new attractors if the potential is steep enough [20], which is very similar to the hyperbolic inflation [8]. This kind of models was firstly constructed to explain the statistical anisotropy in the CMB [21, 22, 23, 24]. The vector field is coupled to gravity through $f^{2}(\phi)F_{\mu\nu}F^{\mu\nu}$. After suitable choices of $f(\phi)$, the dynamics of background vector fields provides an effective potential for the slow-roll scalar field.
@@
$\mathcal{P}_{\zeta}=\mathcal{P}_{\zeta}^{(0)}\left[1+\Theta(\theta)IN^{2}\right],$ (1.1)
where $N\equiv\ln{(-k\tau_{e})}$ represents the e-folding number of observed mode elapsed since horizon-crossing and $I=2\epsilon(\rho_{A}/V)$ is the fraction kinetic energy of the vector field. For one U(1) vector field, $\Theta(\theta)=24\sin^{2}\theta$, where $\theta$ is the angle between mode and direction of the vector field [22, 25]. The term $IN^{2}$ accountable for the amplitude of statistical anisotropy and should be rigorously constrained by CMB observations to $I<10^{-7}$ for a sufficiently lengthy inflationary period $N\sim 60$ [22, 25].The exceedingly small parameter implies the presence of a very small but non-zero vector field, potentially giving rise to fine-tuning challenges [26, 27]. Studies have revealed that employing multiple vector fields can mitigate the anisotropy, eventually converging to an isotropic attractor [24]. For the isotropic configuration of a triad of vector fields, $\Theta(\theta)=16$, relaxing the constraint on $I$ [28]. However, it still needs to be constrained to $I<10^{-5}$due to the tilt of the power spectrum caused by the e-folding number $N$. This is a cumulative effect, i.e., the curvature is sourced by non-zero vector fields and evolves on super-horizon scales.
However, this does not imply that results for large $I$ are undesirable. If we require vector fields to handle a sufficiently large field displacement, indicating that $I$ is not a small parameter, it appears that the second term in (1.1) becomes significant and violates the adiabatic and scale-invariant initial conditions. However, this result lacks reliability because the perturbative method breaks down in this regime. It has been demonstrated that for $I\gtrsim 0.2$, the calculation by using in-in formalism starts to deviate from the numerical results and it is difficult to track the dynamics near and outside the horizon analytically [29, 30]. Numerical findings indicate that the power spectrum does not actually evolve on super-horizon scales [30, 20]. In this paper, we revisit this dilaton-gauge coupling model in comoving gauge and thoroughly explore the $I\gg 1$ case with the assistance of effective field theory (EFT). We will show, the entropy perturbation is super heavy in this regime and the theory can be described by a low-energy EFT of a massless adiabatic perturbation. $IN^{2}$ does not occur in this regime; instead, it is replaced by an exponentially enhanced power spectrum, dependent solely on the parameter $I$. This enhancement also provides a mechanism for generating primordial black holes.
@@ 2 Inflation with vector fields
We are considering the spatially flat Friedmann-Lemaitre-Robertson-Walker (FLRW) universe, which is characterized by the scale factor $a(t)$ and the Hubble variable is $H(t)=\dot{a}/a$. The Arnowitt-Deser-Misner(ADM) formalism of the metric is [31]
@@
where $N$ is the lapse function, $N^{i}$ is the shift vector and $g_{ij}$ is the metric of hypersurface. We consider the model of inflaton and $U(1)$ gauge fields, which is called dilaton-gauge inflation in this paper
$S=\int d^{4}x\sqrt{-{g}}\left[-\frac{1}{2}\partial^{\mu}\phi\partial_{\mu}\phi-\frac{1}{4}f_{ab}(\phi)F^{a}_{\ \mu\nu}F^{b\mu\nu}\right],$ (2.2)
where $f_{ab}$ is the kinetic coupling and only depends on the inflaton.
@@ 2.1 Background
$\ddot{\phi}+3H\dot{\phi}+V_{\phi}-3f_{\phi}f\frac{\dot{\mathbb{A}}^{2}}{a^{2}}=0,\ \ \ \ \ \ \ \ddot{\mathbb{A}}+\left(2\frac{\dot{f}}{f}+H\right)\dot{\mathbb{A}}=0.$ (2.6)
In this paper we assume that $\dot{\phi}<0$ and $\dot{\mathbb{A}}>0$. We also define a useful variable, the ratio of the kinetic energy of $U(1)$ fields and dilaton [^1] The parameter $I$ in [28] is related to $h$ in our paper as $I\simeq 2h^{2}$.
@@
$V_{\phi}=M_{\text{pl}}H^{2}\left(6\epsilon+2\epsilon_{\phi}\eta_{\phi}+\frac{3}{2}\epsilon_{A}\eta_{A}-3\epsilon\epsilon_{A}\right).$ (2.8)
@@ 2.2 Quadratic action in comoving gauge
$N=1+\alpha,\ \ \ \ N_{i}=\partial_{i}\beta,\ \ \ \ g_{ij}=a^{2}(t)e^{2\zeta}\delta_{ij},\ \ \ \ \phi=\phi(t)+\delta\phi.$ (2.11)
@@
where we have discard slow-roll suppressed terms in the slow-low limit and also assumed $|V_{\phi\phi}|/H^{2}\ll 1$. This is the decoupling limit $\epsilon\to 0$ in the EFT. Then $\epsilon_{\phi}$ and $\epsilon_{A}$ are also vanishing because they are positive quantities.
@@ 2.3 Evolution of the adiabatic mode
These constant solutions exist for any value of $h$. And it’s obvious that these mode solutions are not the correction $IN^{2}$ in (1.1). We next explore the Lagrangian for small $h$.
@@
where $N_{c}$ is the time of horizon-crossing and $N_{k}\equiv N_{e}-N_{c}$ is the e-folding number elapsed since horizon-crossing. For small $h$ this is a small correction to the adiabatic mode. The zero-th order entropy perturbation is decoupled with the adiabatic one and massless in the slow-roll limit. Hence the zeroth-order power spectrum of entropy perturbation is simple $\mathcal{P}_{\mathcal{F}}^{(0)}=H^{2}/(4\pi^{2})$. Then the leading-order correction to the power spectrum of curvature perturbation can be estimated to
@@
which is consistent with the result calculated through in-in formalism [28]. These correction is proportional to square of e-folding number $N\sim 60$, which is a large number for inflation. Hence for anisotropic configuration of vector fields, there is very strong constrain from CMB on the parameter $h$. The extreme small $h$ may lead to fine-tuning problem [26, 27]. However, these modes is heavy when $h$ is large enough hence dilute rapidly during inflation. To see this, we next explore the large $h$ case.
@@ 3 EFT under massive entropy mode
$\mathcal{L}_{\text{EFT}}=\sqrt{-g}\left[\frac{M_{\text{pl}}^{2}}{2}R+M_{\text{pl}}^{2}\dot{H}g^{00}-M_{\text{pl}}^{2}\left(3H^{2}+\dot{H}\right)+\cdots\right]+\mathcal{L}_{\text{n-u}},$ (3.1)
$\mathcal{L}_{\text{n-u}}=\sqrt{-g}\sum_{n=2}\frac{M_{n}^{4}}{n!}\left[(1+g^{00})\frac{M^{2}}{M^{2}-\tilde{\partial}^{2}}\right]^{n-1}(1+g^{00})+\cdots,$ (3.2)
$g^{00}=\frac{\partial x^{0}}{\partial\tilde{x}^{\mu}}\frac{\partial x^{0}}{\partial\tilde{x}^{\nu}}\tilde{g}^{\mu\nu}=\left(1+\dot{\pi}\right)^{2}\tilde{g}^{00}+2\left(1+\dot{\pi}\right)\partial_{i}\pi\tilde{g}^{0i}+\tilde{g}^{ij}\partial_{i}\pi\partial_{j}\pi$ (3.3)
In the low-energy scale $\omega\sim p\ll M$, we can write down the action in the re-parameterization of time. Dropping the tildes for simplicity the Lagrangian is given by
$\mathcal{L}_{\text{EFT}}=$ $\sqrt{-g}\Bigg{\{}\frac{M_{\text{pl}}^{2}}{2}R-M_{\text{pl}}^{2}\left[3H^{2}(t+\pi)+\dot{H}(t+\pi)\right]$
$+M_{\text{pl}}^{2}\dot{H}(t+\pi)\left[\left(1+\dot{\pi}\right)^{2}g^{00}+2\left(1+\dot{\pi}\right)\partial_{i}\pi g^{0i}+\partial^{i}\pi\partial_{i}\pi\right]+$
$+\frac{M_{2}^{4}}{2}\left[1+(1+\dot{\pi})^{2}g^{00}+2(1+\dot{\pi})\partial_{i}\pi g^{0i}+\partial^{i}\pi\partial_{i}\pi)\right]^{2}+\cdots\Bigg{\}}$ (3.4)
Expanding the action up to second order of fluctuations, the constrain of the lapse function is solved as $\alpha=\epsilon H\pi$ [41]. In the dilaton-gauge theory, from the quadratic action in the spatial flat gauge (A), constrain $\alpha$ is given by $\alpha\propto Q_{\sigma}$, which means that the unitary gauge is the vanishing adiabatic fluctuation at linear order.
$\mathcal{L}_{\text{EFT}}=a^{3}M_{\text{pl}}^{2}H^{2}\frac{\epsilon}{c_{s}^{2}}\left[\dot{\pi}^{2}-\frac{c_{s}^{2}}{a^{2}}(\partial\mathcal{\pi})^{2}\right],$ (3.5)
where the speed of sound is defined by
$\frac{1}{c_{s}^{2}}\equiv 1-\frac{2M_{2}^{4}}{M_{\text{pl}}^{2}\dot{H}}.$ (3.6)
The modification of the speed of sound is provided by the higher operator $M_{2}$ and we will show soon that this can be obtained by integrating out the heavy fluctuation at low energy in the dilaton-gauge model.
@@ 3.1 Modified dispersion relation
The modification of the dispersion relation of fluctuations is provided by the interaction with $\dot{\zeta}{\mathcal{F}}$. The heavy entropy fluctuation implies the hierarchy of the two frequencies $\omega_{-}\ll\omega_{+}$, which requires $4(1-c_{s}^{2})p^{2}\ll m_{s}^{2}c_{s}^{-2}$. Hence we have a cut-off scale of EFT (3.5) on momentum $p_{\text{UV}}$ and the EFT is only valid below this scale. The UV scale is given by
@@ 3.2 Integrating out the heavy $\mathcal{F}$
We have noted that the speed of sound is $|c_{s}|\sim\mathcal{O}(1)$ for large $h$ hence the non-linear dispersion regime is narrower, and we can directly investigate the linear dispersion regime. In the heavy limit $|m_{s}^{2}|\gg\omega^{2}$ we can expand the operator
@@ 4 Primordial black holes
The remained adiabatic mode experiences a exponential growth before horizon-crossing due to the imaginary speed of sound. Therefore the power spectrum is enhanced exponentially. If vector fields switch on only at late time, The power spectrum is enhanced on small scales so the formations of PBHs in radiation era is possible [47, 48]. And these PBHs can be one of candidates of the dark matter [49, 50, 51] and supermassive black holes in our universe [52].
@@ 4.1 Board case
For this case, we have enough e-folding of $h(N)$ for exponential growth of the modes. The EFT is valid at time $N_{k}$ satisfying
@@
$\Delta_{\zeta}(k)\equiv\frac{\mathcal{P}_{\zeta}(k)}{\mathcal{P}_{\zeta}^{(0)}(k)}\simeq e^{2\mathcal{I}(h)}\big{|}_{N_{k}},$ (4.4)
where the exponential factor is given by (see Appendix B for derivation)
@@ 4.2 Sharp case
Figure 4: The numerical and analytical results (C) of power spectrum for the top-hat case, where $\delta N=0.4$ and $h_{0}=24$. The envelope (4.14) is shown as orange dashed curve.
@@
During the short region-II, where the scale factor is nearly invariant, we have $a=a(N_{f})=k_{f}/H$ hence $p=(k/k_{f})H$ in the above equations. We have eight unknown coefficients $A_{\pm}^{\zeta}$, $B_{\pm}^{\zeta}$, $C_{\pm}^{\zeta}$ and $D_{\pm}^{\zeta}$ in the region-II and we have eight boundary conditions at $t_{1}$ hence all these coefficients can be completely solved. Then at $t_{2}$ we also have eight boundary condictions and the coefficients of (C) can be solved to (see Appendiex C)
@@
where $\kappa\equiv k/k_{f}$, $\delta N\equiv H(t_{2}-t_{1})$ and $S_{\pm}\equiv\omega_{\pm}/H$. We show the agreement between numerical and completed analytical results (see appendix C) in Figure 4. The power spectrum has exponential enhancement near $k_{f}$ and also rapidly oscillate during these scales. We found $E_{\zeta}$ and $F_{\zeta}$ are dominated so the power spectrum can be simply written as $\Delta_{\zeta}\simeq\Delta_{\text{env}}\times\left(\text{rapid oscillation}\right)$, where the envelope is
@@
We have eight unknown coefficients $c_{0}^{i}$, $c_{1}^{i}$, $c_{+}^{i}$ and $c_{-}^{i}$ ($i=\zeta,\mathcal{F}$) in the region-II and eight boundary conditions at $t_{1}$ hence all these coefficients can be completely solved. Then at $t_{2}$ the coefficients of (C) can also be solved [^2] We found that this computation breaks when $x\lesssim 10^{-5}$.. The results are extremely tedious, but we can simple them under some approximations, see appendix C. The results should be real because the perturbations become classical on super-horizon. So we can only consider the real parts

### v2

@@ Abstract
We revisit inflation coupled with vector fields employing kinetic coupling in the comoving gauge. It is known that there is a cumulative effect $\tilde{I}N^{2}$ on curvature power spectrum. For a large number of e-foldings $N$, this contribution is so significant that it could violate observational constraints when the ratio of kinetic energy between vector fields and inflaton $\tilde{I}$ is not extremely small. In this paper, we explore the regime where $\tilde{I}\gg 1$, a realm that has not been extensively explored due to the limitations of perturbative methods. We found that the entropy perturbation becomes heavy in this regime and the cumulative effect decays away on super-horizon scales. Consequently, the power spectrum retains its scale invariance in the decoupling limit. By straightforwardly integrating out the heavy modes near horizon-crossing, we derive a low-energy effective field theory describing a massless adiabatic perturbation with an imaginary speed of sound $c_{s}^{2}=-1/3$. Namely, the inflation with vector fields presents a potential mechanism for generating primordial black holes.
@@ 1 Introduction
The study of introducing extra fields during inflation has attracted attention to evade the issues of single-field inflation. In particular, the study of multi-scalar fields exhibiting strongly non-geodesic motion has been extensively examined in recent years [8, 9, 10, 11, 12, 13, 14]. In contrast, our focus is on U(1) gauge fields, commonly known as vector fields, as additional fields during inflation. This choice stems not only from the ubiquity of gauge fields as essential components of the standard model (SM) and beyond but also from their rich phenomenological implications in cosmology [15, 16, 17, 18, 19]. The breaking of the conformal coupling between vector fields and gravity is necessary to generate sufficiently strong vector fields [19]. Analogous to the non-geodesic motion of scalar fields, vector fields should exhibit background dynamics and participate alongside scalar fields in driving inflation. The inflationary phase undergoes a transition to new attractors if the potential is steep enough [20], which is very similar to the hyperbolic inflation [8]. This kind of models was firstly constructed to explain the statistical anisotropy in the CMB [21, 22, 23, 24] (or read the review [25]). The vector field is coupled to gravity through $f^{2}(\phi)F_{\mu\nu}F^{\mu\nu}$. After suitable choices of $f(\phi)$, the dynamics of background vector fields provides an effective potential for the slow-roll scalar field.
@@
$\mathcal{P}_{\zeta}=\mathcal{P}_{\zeta}^{(0)}\left[1+\Theta(\theta)IN_{k}^{2}\right],\ \ \ \ \ \ (I\ll 1)$ (1.1)
where $N_{k}\equiv\ln{(-k\tau_{e})}$ represents the e-folding number of observed mode elapsed since horizon-crossing, $\tau_{e}$ is time of end of the inflation. $I$ is the parameter related to specific models. For one U(1) vector field, $\Theta(\theta)=24\sin^{2}\theta$, where $\theta$ is the angle between mode and direction of the vector field [22, 26]. The term $IN_{k}^{2}$ accountable for the amplitude of statistical anisotropy and should be rigorously constrained by CMB observations to $I<10^{-7}$ for a sufficiently lengthy inflationary period $N_{k}\sim 60$ [22, 26].The exceedingly small parameter implies the presence of a very small but non-zero vector field, potentially giving rise to fine-tuning challenges [27, 28]. Studies have revealed that employing multiple vector fields can mitigate the anisotropy, eventually converging to an isotropic attractor [24]. For the isotropic configuration of a triad of vector fields, $\Theta(\theta)=16$, relaxing the constraint on $I$ [29]. However, it still needs to be constrained to $I<10^{-5}$ due to the tilt of the power spectrum caused by the e-folding number $N$. This is a cumulative effect, i.e., the curvature is sourced by non-zero vector fields and evolves on super-horizon scales.
However, this does not imply that results for $I\sim\mathcal{O}(1)$ are undesirable. If we require vector fields to handle a sufficiently large field displacement, indicating that $I$ is not a small parameter, it appears that the second term in (1.1) becomes significant and violates the adiabatic and scale-invariant initial conditions. However, this result lacks reliability because the perturbative method breaks down in this regime. It has been demonstrated that for $I\gtrsim 0.2$, the calculation by using in-in formalism starts to deviate from the numerical results and it is difficult to track the dynamics near and outside the horizon analytically [30, 31]. Numerical findings indicate that the power spectrum does not actually evolve on super-horizon scales [31, 20]. In this paper, we revisit this dilaton-gauge coupling model in comoving gauge and thoroughly explore the significant $I$ case with the assistance of effective field theory (EFT). We will show, the entropy perturbation is super heavy in this regime and the theory can be described by a low-energy EFT of a massless adiabatic perturbation. $IN_{k}^{2}$ does not occur in this regime; instead, it is replaced by an exponentially enhanced power spectrum, dependent solely on one parameter. This enhancement also provides a mechanism for generating primordial black holes.
@@ 2 Inflation with vector fields
We are considering the spatially flat Friedmann-Lemaitre-Robertson-Walker (FLRW) universe, which is characterized by the scale factor $a(t)$ and the Hubble variable is $H(t)=\dot{a}/a$, where ˙$\dot{}\equiv d/dt$ is cosmological time derivative. The Arnowitt-Deser-Misner(ADM) formalism of the metric is [32]
@@
where $N$ is the lapse function and we choose the cosmological time $N=1$ for background, $N^{i}$ is the shift vector and $g_{ij}$ is the metric of hypersurface. We consider the model of inflaton and $U(1)$ gauge fields, which is called dilaton-gauge inflation in this paper
$S=\int d^{4}x\sqrt{-{g}}\left[-\frac{1}{2}\partial^{\mu}\phi\partial_{\mu}\phi-V(\phi)-\frac{1}{4}f_{ab}(\phi)F^{a}_{\ \mu\nu}F^{b\mu\nu}\right],$ (2.2)
where $g$ is the determinant of metric, $V(\phi)$ is the inflationary potential and $f_{ab}$ is the kinetic coupling and only depends on the inflaton. The strength of vector is $F^{a}_{\ \mu\nu}=\partial_{\mu}A^{a}_{\ \nu}-\partial_{\nu}A^{a}_{\ \mu}$.
@@ 2.1 Background
$\ddot{\phi}+3H\dot{\phi}+V_{\phi}-3f_{\phi}f\frac{\dot{\mathbb{A}}^{2}}{a^{2}}=0,\ \ \ \ \ \ \ \ddot{\mathbb{A}}+\left(2\frac{\dot{f}}{f}+H\right)\dot{\mathbb{A}}=0,$ (2.6)
where $V_{\phi}\equiv\partial V/\partial\phi$ and $f_{\phi}\equiv\partial f/\partial\phi$. In this paper we assume that $\dot{\phi}<0$ and $\dot{\mathbb{A}}>0$. We also define a useful variable, the ratio of the kinetic energy of $U(1)$ fields and dilaton [^1] The parameter $I$ in [29] or in (1.1) is related to $h$ in our paper as $I\simeq 2h^{2}$ for $h\ll 1$.
@@
$\sqrt{2\epsilon_{\phi}}V_{\phi}=M_{\text{pl}}H^{2}\left(6\epsilon+2\epsilon_{\phi}\eta_{\phi}+\frac{3}{2}\epsilon_{A}\eta_{A}-3\epsilon\epsilon_{A}\right).$ (2.8)
@@ 2.2 Quadratic action in comoving gauge
$N=1+\alpha,\ \ \ \ N_{i}=\partial_{i}\beta,\ \ \ \ g_{ij}=a^{2}(t)\left(e^{2\zeta}\delta_{ij}+2\partial_{i}\partial_{j}E\right),\ \ \ \ \phi=\phi(t)+\delta\phi.$ (2.11)
@@
where we have discarded slow-roll suppressed terms in the slow-low limit and also assumed $|V_{\phi\phi}|/H^{2}\ll 1$. This is the decoupling limit $\epsilon\to 0$ in the EFT. Then $\epsilon_{\phi}$ and $\epsilon_{A}$ are also vanishing because they are positive quantities.
@@ 2.3 Evolution of the adiabatic mode
These constant solutions exist for any value of $h$. And it’s obvious that these mode solutions are not the correction $IN_{k}^{2}$ in (1.1). We next explore the Lagrangian for small $h$.
@@
where $N_{c}$ is the time of horizon-crossing, $N_{e}$ is the time of end of inflation and $N_{k}\equiv N_{e}-N_{c}$ is the e-folding number elapsed since horizon-crossing. For small $h$ this is a small correction to the adiabatic mode. The zero-th order entropy perturbation is decoupled with the adiabatic one and massless in the slow-roll limit. Hence the zeroth-order power spectrum of entropy perturbation is simple $\mathcal{P}_{\mathcal{F}}^{(0)}=H^{2}/(4\pi^{2})$. Then the leading-order correction to the power spectrum of curvature perturbation can be estimated to
@@
which is consistent with the result calculated through in-in formalism [29] and reproduces the result (1.1). These correction is proportional to square of e-folding number $N_{k}\sim 60$, which is a large number for inflation. Hence for anisotropic configuration of vector fields, there is very strong constrain from CMB on the parameter $h$. The extreme small $h$ may lead to fine-tuning problem [27, 28]. However, these modes is heavy when $h$ is large enough hence dilute rapidly during inflation. To see this, we next explore the large $h$ case.
@@ 3.1 Modified dispersion relation
The modification of the dispersion relation of fluctuations is provided by the interaction with $\dot{\zeta}{\mathcal{F}}$. The heavy entropy fluctuation implies the hierarchy of the two frequencies $\omega_{-}\ll\omega_{+}$, which requires $4(1-c_{s}^{2})p^{2}\ll m_{s}^{2}c_{s}^{-2}$. Hence we have a cut-off scale of EFT (B.8) on momentum $p_{\text{UV}}$ and the EFT is only valid below this scale. The UV scale is given by
@@ 3.2 Integrating out the heavy $\mathcal{F}$
We have noted that the speed of sound is $|c_{s}|\sim\mathcal{O}(1)$ for large $h$ hence the non-linear dispersion regime is narrower, and we can directly investigate the linear dispersion regime (see Appendix B for reviews). In the heavy limit $|m_{s}^{2}|\gg\omega^{2}$ we can expand the operator
@@ 4 Primordial black holes
The remained adiabatic mode experiences a exponential growth before horizon-crossing due to the imaginary speed of sound. Therefore the power spectrum is enhanced exponentially. If vector fields switch on only at late time, The power spectrum is enhanced on small scales so the formations of PBHs in radiation era is possible [48, 49]. And these PBHs can be one of candidates of the dark matter [50, 51, 52] and supermassive black holes in our universe [53]. The enhancement of the power spectrum due to the heavy fields is studied recently in multiple scalar-field models, and is used to discuss the PBHs formation [56, 54, 55]. In previous section, we have realized that it also can be achieved by using vector fields.
@@ 4.1 Board case
For this case, we have enough e-folding of $h(N)$ for exponential growth of the modes. The EFT is valid at time $N_{v}$ satisfying
@@
$\Delta_{\zeta}(k)\equiv\frac{\mathcal{P}_{\zeta}(k)}{\mathcal{P}_{\zeta}^{(0)}(k)}\simeq e^{2\mathcal{I}(h)}\big{|}_{N_{v}},$ (4.4)
where the exponential factor is given by (see Appendix C for derivation [^2] Although this analytical result fixes well with the numerical one. This derivation ignores the Hubble friction in the equations of motion. After completing this paper, [58] proposed a rigorous derivation when take into account the Hubble fraction terms.)
@@ 4.2 Sharp case
Figure 4: The numerical and analytical results (D) of power spectrum for the top-hat case, where $\delta N=0.4$ and $h_{0}=24$. The envelope (4.14) is shown as orange dashed curve.
@@
During the short region-II, where the scale factor is nearly invariant, we have $a=a(N_{f})=k_{f}/H$ hence $p=(k/k_{f})H$ in the above equations. We have eight unknown coefficients $A_{\pm}^{\zeta}$, $B_{\pm}^{\zeta}$, $C_{\pm}^{\zeta}$ and $D_{\pm}^{\zeta}$ in the region-II and we have eight boundary conditions at $t_{1}$ hence all these coefficients can be completely solved. Then at $t_{2}$ we also have eight boundary condictions and the coefficients of (4.2) can be solved to (see Appendiex D)
@@
where $\kappa\equiv k/k_{f}$, $\delta N\equiv H(t_{2}-t_{1})$ and $S_{\pm}\equiv\omega_{\pm}/H$. We show the agreement between numerical and completed analytical results (see appendix D) in Figure 4. The power spectrum has exponential enhancement near $k_{f}$ and also rapidly oscillate during these scales. We found $E_{\zeta}$ and $F_{\zeta}$ are dominated so the power spectrum can be simply written as $\Delta_{\zeta}\simeq\Delta_{\text{env}}\times\left(\text{rapid oscillation}\right)$, where the envelope is
@@
We have eight unknown coefficients $c_{0}^{i}$, $c_{1}^{i}$, $c_{+}^{i}$ and $c_{-}^{i}$ ($i=\zeta,\mathcal{F}$) in the region-II and eight boundary conditions at $t_{1}$ hence all these coefficients can be completely solved. Then at $t_{2}$ the coefficients of (4.2) can also be solved [^3] We found that this computation breaks when $x\lesssim 10^{-5}$.. The results are extremely tedious, but we can simple them under some approximations, see appendix D. The results should be real because the perturbations become classical on super-horizon. So we can only consider the real parts